# Stochastic Parameter Decomposition

### v1

@@ A.2 Avoiding dead gradients with leaky hard sigmoids
The flat regions in a hard sigmoid function can lead to dead gradients for inputs below 00 or above $1$. To avoid this, we use leaky hard sigmoids instead, which introduce a small non-zero slope below 00 or above $1$. Specifically, we use lower-leaky hard sigmoids $\sigma_{H,\text{lower}}(x)$ with slope $0.01$ below 00 for the gates used in the forward pass for the $\mathcal{L}_{\text{stochastic-recon}}$ and $\mathcal{L}_{\text{stochastic-recon-layerwise}}$ losses. And we use upper-leaky hard sigmoids $\sigma_{H,\text{upper}}(x)$ with slope $0.01$ above $1$ in the causal importance loss $\mathcal{L}_{\text{importance-minimality}}$ :
$\sigma_{H,\text{lower}}(x):=\begin{cases}0.01x&x\leq 0\\ x&0\leq x\leq 1\\ 1&1\leq x\end{cases}$ (14)
$\sigma_{H,\text{upper}}(x):=\begin{cases}0&x\leq 0\\ x&0\leq x\leq 1\\ 1+0.01(x-1)&1\leq x\end{cases}$
@@ A.4.2 Toy models of Compressed computation and Cross-Layer Distributed Representation
All models trained using AdamW with weight decay 0.01, max learning rate $3\times 10^{-3}$ with cosine decay, batch size 2048. The dataset uses input features sampled from $[-1,1]$ uniformly with probability 0.01, and 0 otherwise.
@@ A.5 Supplementary figures
Figure 14: Toy Model of Distributed Representations (Two Layers): Similarity between target model weights and SPD model components for the first $10$ input feature dimensions in a $2$-layer residual MLP. Top: Neuron contributions measured by $(W_{E}W_{\text{IN}})\odot(W_{\text{OUT}}W_{U})$, where $\odot$ is an element-wise product and $W_{\text{IN}}$ and $W_{\text{OUT}}$ are the MLP input and output matrices in both layers concatenated together. Bottom: Neuron contributions for the learned parameter components, measured by $\max_{m}[({W_{U}}_{[i,:]}U^{\text{OUT}}V^{\text{OUT}})\odot(U^{\text{IN}}_{[:,m]}V^{\text{IN}}_{[m,:]}{W_{E}}_{[:,i]})]$ for each feature index $i\in[0,9]$. The neurons are numbered based on their raw position in the network, with neurons 00 to $24$ in the first layer and neurons $25$ to $49$ in the second layer.
Figure 15: Toy Model of Distributed Representations (Three Layers): Similarity between target model weights and SPD model components for the first $10$ input feature dimensions in a $3$-layer residual MLP. Top: Neuron contributions measured by $(W_{E}W_{\text{IN}})\odot(W_{\text{OUT}}W_{U})$ where $\odot$ is an element-wise product and $W_{\text{IN}}$ and $W_{\text{OUT}}$ are the MLP input and output matrices in all three layers concatenated together. Bottom: Neuron contributions for the learned parameter components, measured by $\max_{m}[({W_{U}}_{[i,:]}U^{\text{OUT}}V^{\text{OUT}})\odot(U^{\text{IN}}_{[:,m]}V^{\text{IN}}_{[m,:]}{W_{E}}_{[:,i]})]$ for each feature index $i\in[0,9]$. The neurons are numbered based on their raw position in the network, with neurons 00 to $16$ in the first layer, $17$ to $33$ in the second layer and neurons $34$ to $50$ in the third layer.

### v2

@@ A.2 Avoiding dead gradients with leaky hard sigmoids
The flat regions in a hard sigmoid function can lead to dead gradients for inputs below 0 or above $1$. To avoid this, we use leaky hard sigmoids instead, which introduce a small non-zero slope below 0 or above $1$. Specifically, we use lower-leaky hard sigmoids $\sigma_{H,\text{lower}}(x)$ with slope $0.01$ below 0 for the gates used in the forward pass for the $\mathcal{L}_{\text{stochastic-recon}}$ and $\mathcal{L}_{\text{stochastic-recon-layerwise}}$ losses. And we use upper-leaky hard sigmoids $\sigma_{H,\text{upper}}(x)$ with slope $0.01$ above $1$ in the causal importance loss $\mathcal{L}_{\text{importance-minimality}}$ :
[…]
@@ A.4.2 Toy models of Compressed computation and Cross-Layer Distributed Representation
All models trained using AdamW with weight decay 0.01, max learning rate $3\times 10^{-3}$ with cosine decay, batch size 2048. The dataset uses input features sampled from [1,1][-1,1] uniformly with probability 0.01, and 0 otherwise.
@@ A.5 Supplementary figures
Figure 14: Toy Model of Distributed Representations (Two Layers): Similarity between target model weights and SPD model components for the first $10$ input feature dimensions in a 22-layer residual MLP. Top: Neuron contributions measured by $(W_{E}W_{\text{IN}})\odot(W_{\text{OUT}}W_{U})$, where $\odot$ is an element-wise product and $W_{\text{IN}}$ and $W_{\text{OUT}}$ are the MLP input and output matrices in both layers concatenated together. Bottom: Neuron contributions for the learned parameter components, measured by $\max_{m}[({W_{U}}_{[i,:]}U^{\text{OUT}}V^{\text{OUT}})\odot(U^{\text{IN}}_{[:,m]}V^{\text{IN}}_{[m,:]}{W_{E}}_{[:,i]})]$ for each feature index $i\in[0,9]$. The neurons are numbered based on their raw position in the network, with neurons 0 to $24$ in the first layer and neurons $25$ to $49$ in the second layer.
Figure 15: Toy Model of Distributed Representations (Three Layers): Similarity between target model weights and SPD model components for the first $10$ input feature dimensions in a 33-layer residual MLP. Top: Neuron contributions measured by $(W_{E}W_{\text{IN}})\odot(W_{\text{OUT}}W_{U})$ where $\odot$ is an element-wise product and $W_{\text{IN}}$ and $W_{\text{OUT}}$ are the MLP input and output matrices in all three layers concatenated together. Bottom: Neuron contributions for the learned parameter components, measured by $\max_{m}[({W_{U}}_{[i,:]}U^{\text{OUT}}V^{\text{OUT}})\odot(U^{\text{IN}}_{[:,m]}V^{\text{IN}}_{[m,:]}{W_{E}}_{[:,i]})]$ for each feature index $i\in[0,9]$. The neurons are numbered based on their raw position in the network, with neurons 0 to $16$ in the first layer, $17$ to $33$ in the second layer and neurons $34$ to $50$ in the third layer.